OpenStudy (anonymous):

\[\huge \int \frac{arctan(x)}{1+x^2} dx \]

OpenStudy (ash2326):

put \(x= tan \theta\) \[dx= \sec^2 \theta\] we get \[\int \frac { \theta \sec^2 \theta}{sec^2 \theta} d\theta\] we have now \[\int \theta d\theta= \frac{{\theta}^2}{2} +c\] substitute \(\theta = \tan ^{-1} x\)

OpenStudy (campbell_st):

let u = arctan(x) then du/dx = 1/(1+x^2) or du = 1/(1+x^2) dx so the problem is now \[\int\limits u du\] \[= 1/2 u^2 + c\] = 1/2 (arctan(x))^2 + c

OpenStudy (anonymous):

campbell is right

OpenStudy (anonymous):

got the same

OpenStudy (campbell_st):

oops it should read \[u^2/2 + c\]